# Callan-Symanzik-like equation in information theory

### v1

@@ 5 Conclusion
The AdS/CFT by corresponding a gravitational theory to a quantum theory could alleviate the perturbation calculation in strongly-coupled systems. Exploring the counterpart quantities in the AdS/CFT make it a strong technique to study the physical theories. Computational complexity as one of these quantities has been considered in holography and provided different proposal. Recently, a proposal known as complexity equals anything has been introduced in which the volume of a generalized surface is corresponded to the complexity of the boundary theory. Studying the complexity of various models through this proposal shows some branches in CGR. Moreover, these branches has been considered as a phase transition in CGR.
Because of the fact that the phase transition appeared in complexity equals anything, this suggests that the generalized surfaces could make the transition. One of the mechanism for the generalization is the Weyl squared tensor. Expanding the Weyl squared tensor in the Fefferman-Graham coordinate makes it to be reduced into three terms: divergence of the energy-momentum tensor, shear viscosity and the Weyl anomaly of the boundary theory. Then, the phase transition and the distance between the jumps is controlled by these three terms.
CGR in this way behaves as a function of the generalization parameter and the locations of the jumps. In other words, these two make the CGR to be changed. Considering the behavior of CGR around the phase transition reveals a scaling behavior and critical exponents. It means that the interpretation of the phase transition our of these jumps makes sense. However, CGR satisfies a Callan-Symanzik-like (CSL) equation in field theory in which a beta function is defined in terms of the generalization parameter and an anomalous dimension in terms of the scaling of CGR, which means that the critical behavior of CGR is universal regardless of the generalization parameter.
CSL equation for CGR suggests that in different scale of energy, the minimum required quantum gates is changed. In other words, the change in the scale of energy could change the speed of the processing of information. This recommends that one way to change the speed of the information processing is changing the scale of energy. Suppose that we prepare a quantum computer with a particular speed of processing. One could instead of working on the quantum gates to improve the the speed of processing, work on the scale of energy, put it another way, with a particular array of quantum gates it is enough to change the scale of energy to have fast processing. An interesting application of CSL would be turning the classical computers into a faster one and even faster than our current quantum computer by changing the scale of energy not by the gates or architecture of gates.
@@ Acknowledgment
Author would like to express his gratitude to Monireh Emami for useful discussions.
@@ Appendix A Weyl squared tensor in FG coordinate
The line element in Fefferman-Graham coordinate is given by:
@@
where $R_{ij}$ and $R$ are the Ricci tensor and the Ricci scalar of the boundary, respectively. The Weyl tensor, the Riemann tensor and the Ricci tensor by definition could be written, respectively:
$\mathcal{C}_{\mu\nu\rho\sigma}=R_{\mu\nu\rho\sigma}-\Big{(}g_{\mu[\rho}P_{\sigma]\nu}-g_{\nu[\rho}P_{\sigma]\mu}\Big{)},$ (A-4)
$\mathcal{C}_{\rho i\rho j}=R_{\rho i\rho j}-\frac{R_{\rho\rho}}{3}g^{(0)}_{ij}+\mathcal{O}(\rho^{4}),$ (A-5)
$R_{\rho i\rho j}=-\frac{\rho^{2}}{l^{2}}\Big{(}<T_{ij}>-\frac{g^{(0)}_{ij}}{3}<T_{k}^{k}>\Big{)}+\mathcal{O}(\rho^{4}),$ (A-6)
$R_{\rho\rho}=-\frac{4}{l^{2}}+\mathcal{O}(\rho^{2}).$ (A-7)
Then, the Weyl squared tensor in terms of the boundary metric would be:
$\mathcal{C}_{\rho i\rho j}\mathcal{C}^{\rho i\rho j}\sim\Big{(}<T_{ij}>-\frac{g^{(0)}_{ij}}{4}<T_{k}^{k}>\Big{)}^{2}+\mathcal{O}(\rho^{6}).$ (A-8)
In addition, for the other component [9]:
$R_{\rho ijk}\sim\nabla_{i}P_{jk}-\nabla_{j}P_{ik},$ (A-9)
$C_{\rho ijk}C^{\rho ijk}\sim\Big{(}\nabla_{i}P_{jk}-\nabla_{j}P_{ik}\Big{)}^{2}+\mathcal{O}(\rho^{4}),$ (A-10)
$\nabla^{i}g^{(4)}_{ij}=\nabla^{i}P_{ik}P_{j}^{k}-P_{ik}\nabla_{j}P^{ik}.$ (A-11)
The leading term would be:
$\nabla^{i}g^{(4)}_{ij}\sim\nabla^{i}<T_{ij}>\sim\nabla^{i}P_{ij},$ (A-12)
$\mathcal{C}_{\rho ijk}\mathcal{C}^{\rho ijk}\sim\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{4})$ (A-13)

### v2

@@ 5 Conclusion
The AdS/CFT correspondence, by relating a gravitational theory to a quantum field theory, can alleviate the difficulties of perturbation theory in strongly-coupled systems. Exploring the counterpart quantities in the AdS/CFT make it a strong technique to study the physical theories. Computational complexity, as one of these quantities, has been extensively studied in holography through several proposals. Recently, a proposal known as complexity equals anything has been introduced where the complexity is identified with the volume of a generalized hypersurface. Studying the complexity of various models through this proposal shows some branches in CGR. Moreover, these branches has been considered as a phase transition in CGR.
Since phase transitions appear in the ”complexity = anything” framework, this suggests that the generalized surfaces could make the transition. One mechanism for the generalization is the Weyl squared tensor. Expanding the Weyl squared tensor in the Fefferman-Graham coordinate makes it to be reduced into two terms: divergence of the energy-momentum tensor and the Weyl squared tensor of the boundary which is only a geometrical expansion. However, the phase transitions and jump separations are governed by the boundary energy-momentum tensor.
CGR in this way behaves as a function of the generalization parameter and the locations of the jumps. In other words, these two parameters control CGR. Considering the behavior of CGR around the phase transition reveals a scaling behavior and critical exponents. This shows that interpreting these jumps as phase transitions is meaningful. However, CGR satisfies a Callan-Symanzik-like (CSL) equation in field theory in which a beta function is defined in terms of the generalization parameter and an anomalous dimension in terms of the scaling of CGR, which means that the critical behavior of CGR is universal regardless of the dimension.
CSL equation for CGR suggests that in different scale of energy, the minimum required quantum gates is changed. In other words, the change in the scale of energy could change the speed of the processing of information. This may suggest that modifying the energy scale could provide a means to control information processing speed. Suppose that we prepare a quantum computer with a particular speed of processing. One could instead of working on the quantum gates to improve the the speed of processing, modify the energy scale to achieve faster processing. As a future work one could use the application of CSL to evaluate how it improves the processing in real quantum computers.
@@ Acknowledgment
The authors would like to express their gratitude to Monireh Emami for useful discussions. In addition, OpenAI’s ChatGPT has been used to enhance the clarity of this manuscript.
@@ Appendix A Weyl squared tensor in FG coordinate
In this section the derivation of (24) is provided. The line element in Fefferman-Graham coordinate is given by:
@@
In $4+1$ dimensions (d=4) the expansion would be:
$g_{ij}(x,\rho)=g^{(0)}_{ij}(x)+\rho^{2}g^{(2)}_{ij}(x)+\rho^{4}\Big(g^{(4)}_{ij}(x)+h^{(4)}_{ij}(x)Log\rho^{2}\Big)+\mathcal{O}(\rho^{6}).$ (A-4)
Then, the Riemann tensor is given by:
$\Gamma_{\rho\rho}^{\rho}=-\frac{1}{\rho},$ (A-5)
$\Gamma_{ij}^{\rho}=-\frac{1}{\rho}g_{ij}+\frac{1}{2}\partial_{\rho}g_{ij}=2\rho g^{(2)}_{ij}+4\rho^{3}(g^{(4)}_{ij}+h^{(4)}_{ij}Log\rho^{2})+$..., (A-6)
$\Gamma_{\rho j}^{i}=-\frac{1}{\rho}\delta_{j}^{i}+\frac{1}{2}g^{ik}\partial_{\rho}g_{kj}$ (A-7)
$R_{ijkl}(G)=R_{ijkl}(g|_{\rho})-\frac{1}{\rho^{2}}(g_{ik}g_{jl}-g_{il}g_{jk})+\frac{1}{4}\partial_{\rho}g_{ik}\partial_{\rho}g_{jl}-\frac{1}{4}\partial_{\rho}g_{il}g_{jk}+$... (A-8)
where $g|_{\rho}=g^{(0)}_{ij}+\rho^{2}g^{(2)}_{ij}$+... . For the other component:
$R_{\rho i\rho j}(G)=-\frac{1}{2}\partial_{\rho}^{2}g_{ij}+\frac{1}{4}g^{kl}\partial_{\rho}g_{ik}\partial_{\rho}g_{jl}=-\frac{1}{\rho^{2}}g^{(0)}_{ij}$+.... (A-9)
The other component goes as:
$R_{\rho ijk}(G)=\frac{1}{2}(\nabla_{j}\partial_{\rho}g_{ik}-\nabla_{k}\partial_{\rho}g_{ij})=\nabla_{j}(\rho g^{(2)}_{ik})-\nabla_{k}(\rho g^{(2)}_{ij})$. (A-10)
where $\nabla$ is with respect to $g|_{\rho}$. The Ricci tensor would be:
$R_{\rho\rho}=-\frac{1}{2}g^{ij}\partial_{\rho}^{2}g_{ij}+\frac{1}{4}g^{ik}g^{jl}\partial_{\rho}g_{ij}\partial_{\rho}g_{kl},$ (A-11)
$R_{\rho i}=\frac{1}{2}g^{jk}(\nabla_{k}\partial_{\rho}g_{ij}-\nabla_{i}\partial_{\rho}g_{jk}),$ (A-12)
$R_{ij}=R_{ij}(g|_{\rho})-\frac{1}{2}\partial_{\rho}^{2}g_{ij}+\frac{1}{4}g^{kl}\partial_{\rho}g_{ik}\partial_{\rho}g_{jl}+...=R_{ij}(g|_{\rho})-4g^{(2)}_{ij}$+.... (A-13)
The Ricci scalar would be:
$R=-\frac{20}{l^{2}}+\rho^{2}R(g^{(0)})+.$... (A-14)
The Weyl tensor is given by:
$\mathcal{C}_{ijkl}(G)=\mathcal{C}_{ijkl}(g^{(0)})+\rho^{2}g^{(2)}+\rho^{4}(g^{(4)}+h^{(4)}Log\rho^{2})+.$... (A-15)
Now, the Weyl squared tenor is computed as:
$\mathcal{C}^{2}=\rho^{4}\mathscr{C}+\rho^{4}Log\rho^{2}\mathcal{A}+\rho^{6}\mathcal{B}+\rho^{8}\Big((\nabla_{i}g^{(4)}_{jk})^{2}+g^{(4)}_{ij}g^{(4)ij}\Big)+$... (A-16)